OpenStudy (anonymous):

The three sides of an equilateral triangle are increased by 20cm, 30cm, and 40cm, respectively. The resulting triangle's perimeter is between twice and three times the perimeter of the original triangle. What can you conclude about the length of the side ofthe original triangle?

OpenStudy (pfenn1):

Can you write an equation for the resulting perimeter of the triangle if you assume x = length of the side of the original equilateral triangle?

OpenStudy (anonymous):

I don't understand?

OpenStudy (pfenn1):

What is the equation for the perimeter of a triangle?

OpenStudy (anonymous):

x+x+x ?

OpenStudy (pfenn1):

Correct. So the perimeter of the new triangle (where the three sides are increased by 20cm, 30cm, and 40cm, respectively) is given by what equation?

OpenStudy (anonymous):

(x+20) + (x+30) + (x+40) ?

OpenStudy (pfenn1):

Correct.

OpenStudy (pfenn1):

The resulting triangle's perimeter is between twice and three times the perimeter of the original triangle. So if P1 is the original perimeter and P2 is the perimeter of the resultant triangle then \[2P _{1}\le P _{2}\le3P _{1}\] Substitute in the values for P1 and P2 in terms of x

OpenStudy (pfenn1):

\[P_1=x+x+x=3x\]\[P_2=(x+20)+(x+30)+(x+40)=3x+90\]

OpenStudy (anonymous):

How do I solve that as an inequality?

OpenStudy (pfenn1):

You can solve them independently. \[6x \le 3x+90 \le 9x\]\[3x+90 \ge 6x\]and \[3x+90 \le 9x\]

OpenStudy (pfenn1):

Do you know how to do this?

OpenStudy (anonymous):

Yes, I do. The answer I got was X is between 15 and 30. Is that correct?

OpenStudy (pfenn1):

Yep. Great!

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (pfenn1):

You are so welcome!